A rocket takes off from Earth's surface, accelerating straight up at 75.2 m/s2. Calculate the normal force (in N) acting on an astronaut of mass 84.9 kg, including his spacesuit. (Assume the rocket's initial motion is parallel to the +y-direction. Indicate the direction with the sign of your answer.)


___N


1
Expert's answer
2021-09-23T08:33:18-0400

"a=75.2m\/s;"

"m=84.9kg;"

"F=mg+ma;"

"F=m(g+a);"

"F=84.9\sdot(9.8+75.2)=7216.5N;"

Answer: "7216.5N."